If the capacitor in has 0.02 m​2 plates that are 680.76 pm (pico meters) away from one another, determine what dielectric material is in between the plates if the computer was connected to a 100 Volt power supply?
1
Expert's answer
2019-04-04T08:53:10-0400

We have


"S = 0.02 (m^2); d = 680.76 (pm); U = 100 (V); E=6.5 (J)"

So, capacity

"C=\frac{\epsilon\epsilon_0 S}{d}"

Energy of capacitor

"E=\frac{CU^2}{2}"

To find dielectric material

"\epsilon=\frac{2Ed}{\epsilon_0 S U^2}"

We have

"\epsilon=\frac{2 \cdot 6.5 \cdot 680.76\cdot 10^{-12}}{8.85\cdot10^{-12}\cdot 0.02\cdot 100^2}=5"

Answer: Maybe Octyl bromide or Pyrex Glass or porcelain (ceramic)